# A subcopula characterization of dependence for the Multivariate Bernoulli Distribution

### v1

@@ Abstract
By applying Sklar’s theorem to the Multivariate Bernoulli Distribution (MBD), it is proposed a framework that decouples the marginal distributions from the dependence structure, providing a clearer understanding of how binary variables interact. Explicit formulas are derived under the MBD using subcopulas to introduce dependence measures for interactions of all orders, not just pairwise. A bayesian inference approach is also applied to estimate the parameters of the MBD, offering practical tools for parameter estimation and dependence analysis in real-world applications. The results obtained contribute to the application of subcopulas of multivariate binary data, with a real data example of comorbidities in COVID-19 patients.
Keywords: Sklar theorem, subcopula, dependence, Bernoulli, binary data.
@@ 1 Introduction
A representation of the MBD that provides an alternative to the traditional log-linear model for binary variables was proposed by [17] using the concept of Kronecker product from matrix calculus, in terms of $12^{n}-$ 1 parameters, where $n$ is the dimension of a random vector with MBD, but where the dependency vector of parameters are, in fact, non standardized central moments. In [11] several flexible methods for simulating random binary sequences with fixed marginal distributions and specified degrees of association between the variables are discussed, but only pairwise dependencies are considered. An algorithm is derived in [7] for generating systems of correlated binary data, but allowing for the specification of just pairwise correlations within each system. A simple method to characterize multivariate Bernoulli variables with given means is investigated in [5] but only pairwise are correlations are discussed.
Copulas, and more generally subcopulas, provide a straightforward approach to describing the dependence structure in multivariate settings by decoupling the marginal distributions from the joint distribution. Introduced by Sklar’s theorem [15], subcopulas extend copula theory to cases where the marginals are not necessarily continuous, making them particularly useful for multivariate discrete data. This paper focuses on leveraging subcopulas to characterize the dependence within the MBD, offering a framework to better understand and quantify dependence of all orders (pariwise, three-wise, four-wise, etc.)
Building upon previous work in copula theory and dependence measures, the present work introduces a formal subcopula-based characterization of the dependence structure within the MBD, and illustrates its usefulness to tackle the compatibility problem of building multivariate models in terms of lower dimension marginals.
@@ 2 Subcopulas and dependence
Sklar’s Theorem [15] proves that for any $n-$dimensional vector of random variables $(X_{1},\ldots,X_{n})$ there exists a functional link $S^{(n)}$ between its probability joint distribution $F_{X_{1},\ldots,X_{n}}(x_{1},\ldots,x_{n})=\mathbb{P}(X_{1}\leq x_{1},\ldots,X_{n}\leq x_{n})$ and its corresponding univariate marginal distributions $F_{X_{i}}(x_{i})=\mathbb{P}(X_{i}\leq x_{i}),$ $i\in\{1,\ldots,n\}:$
@@
for any $(x_{1},\ldots,x_{n})\in\overline{\mathbb{R}}^{n},$ and where the function $S^{(n)}:D_{1}\times\cdots\times D_{n}\rightarrow[0,1]$ is unique, with $D_{i}=\mbox{Ran}\,F_{X_{i}}$ for $i\in\{1,\ldots,n\},$ which implies $\{0,1\}\subseteq D_{i}\subseteq[0,1].$ Such link $S^{(n)}$ is called a subcopula function, and in the particular case where every $D_{i}=[0,1]$ then it is called a copula function, which would be the case when all the random variables are continuous. If at least one of the random variables is not continuous then the domain of the underlying subcopula $S^{(n)}$ is a proper subset of the unit hypercube $[0,1]^{n}.$ A formal definition for subcopula functions and their properties are discussed in detail in [13] and [12].
Recalling that the random variables $X_{1},\ldots,X_{n}$ are independent if and only if their probability joint distribution is equal to the product of its univariate marginals, then as an immediate consequence of (1) the unique subcopula function under independence is:
@@
Moreover, as a consequence of the Fréchet-Hoeffding bounds for joint distributions (see [6] and [8]) and applying (1) we may obtain bounds for any subcopula $S^{(n)}:$
@@
From any $n-$dimensional subcopula $S^{(n)}(u_{1},\ldots,u_{n})$ and $2\leq k<n$ it is possible to obtain $\binom{n}{k}$ $k-$dimensional marginal subcopulas $S^{(k)}(v_{1},\ldots,v_{k})$ defined in terms of $S^{(n)}(u_{1},\ldots,u_{n})$ where $k$ of its entries are equal to exactly one of each $\{v_{1},\ldots,v_{k}\}$ and the remaining $n-k$ are equal to $1$
Since the univariate marginal distributions $F_{X_{i}}(x_{i})=\mathbb{P}(X_{i}\leq x_{i})$ have no information about how each random variable $X_{i}$ interacts with others, then as a consequence of (1) all the information about the dependence among the random variables $X_{1},\ldots,X_{n}$ is contained in their unique underlying subcopula $S^{(n)},$ and therefore any attempt to measure degrees or intensity of dependence should extract information from $S^{(n)}.$ For the particular case of bivariate copulas (when both random variables are continuous) several copula-based measures have been studied, such as the concordance measures by Spearman [16] or Kendall [10], or dependence measures by Schweizer-Wolff [14] or Hoeffding [8]. For the case of proper subcopulas (that are not copulas) a subcopula dependence measure has been proposed by Erdely [3].
@@ 3 Bivariate Bernoulli
Accordingly to (1) the domain of the underlying subcopula in this case is the set $\mbox{Dom}\,S^{(2)}=\{0,\theta_{1},1\}\times\{0,\theta_{2},1\}.$ From the formal definition of bivariate subcopula (see [12], for example) it always satisfies $S^{(2)}(u,0)=0=S^{(2)}(0,v),$ $S^{(2)}(u,1)=u,$ and $S^{(2)}(1,v)=v,$ and therefore:
$S^{(2)}(u,v)=\begin{cases}\,1&\mbox{ if }(u,v)=(1,1),\\ \,\theta_{1}&\mbox{ if }(u,v)=(\theta_{1},1),\\ \,\theta_{2}&\mbox{ if }(u,v)=(1,\theta_{2}),\\ \,\theta_{12}&\mbox{ if }(u,v)=(\theta_{1},\theta_{2}),\\ \,0&\mbox{ elsewhere,}\end{cases}$ (4)
@@
Formulas (8) allow to fully characterize the family $\mathcal{B}_{2}(\theta_{1},\theta_{2},\theta_{12})$ of all bivariate Bernoulli distributions in terms of the marginal parameters $10<\theta_{1}<$ 1 and $0<\theta_{2}<1,$ and a bivariate dependence parameter $\theta_{12}$ that must satisfy $\max\{\theta_{1}+\theta_{2}-1,0\}\leq\theta_{12}\leq\min\{\theta_{1},\theta_{2}\}.$ The subcopula based dependence measure proposed by [3] for $\mathcal{B}_{2}(\theta_{1},\theta_{2},\theta_{12})$ becomes:
@@
$(p_{00},p_{01},p_{10},p_{11})=(\theta_{1}\theta_{2},\theta_{1}(1-\theta_{2}),(1-\theta_{1})\theta_{2},(1-\theta_{1})(1-\theta_{2}))\,,$ (10)
@@
$(p_{00},p_{01},p_{10},p_{11})=(\theta_{1},0,\theta_{2}-\theta_{1},1-\theta_{2})\,.$ (11)
@@
$(p_{00},p_{01},p_{10},p_{11})=\begin{cases}\,(\theta_{1}+\theta_{2}-1,1-\theta_{2},1-\theta_{1},0)&\mbox{if }\theta_{1}+\theta_{2}>1\,,\\ \,(0,\theta_{1},\theta_{2},1-\theta_{1}-\theta_{2})&\mbox{if }\theta_{1}+\theta_{2}\leq 1\,,\end{cases}$ (13)
@@
In summary, formulas (8) allow to define the family $\mathcal{B}_{2}(\theta_{1},\theta_{2},\theta_{12})$ of all bivariate distributions with Bernoulli univariate marginals, which is in fact also known as a Fréchet class of the type $\mathcal{F}(F_{1},F_{2}),$ see for example [9]. The bivariate joint probability mass function $p_{ij}=\mathbb{P}(X_{1}=i,X_{2}=j)$ for $i$ and $j$ in $\{0,1\}$ can be specified in two ways:
@@
    choosing freely the two marginal parameters $10<\theta_{1}<$ 1 and $10<\theta_{2}<$ 1 and then choosing $\theta_{12}$ such that (5) holds, or
@@
    choosing freely the dependence parameter $10<\theta_{12}<$ 1 and then choosing the marginal parameters $10<\theta_{1}<$ 1 and $10<\theta_{2}<$ 1 such that (5) holds. Equivalently, instead of choosing the dependence parameter value, we may choose a given value of the dependence measure defined by [3] and then obtain the value for $\theta_{12}$ from (9).
Behind the above cases a) and b) is what is known as a compatibility problem, in this particular case for the Fréchet class $\mathcal{F}(F_{1},F_{2}):$ eventhough the three parameters are the probabilities $\theta_{1}=\mathbb{P}(X_{1}=0),$ $\theta_{2}=\mathbb{P}(X_{2}=0),$ and $\theta_{12}=\mathbb{P}(X_{1}=0,X_{2}=0),$ not all values in the $[0,1]$ interval are simultaneously admissible (or compatible) for such parameters, since (5) must hold, see Figure 1.
@@ 4 Trivariate Bernoulli
Let $(X_{1},X_{2},X_{3})$ be a vector of Bernoulli random variables with (univariate marginal) parameters $1-\theta_{r}$ where $10<\theta_{r}<$ 1 as explained in the previous section, but now $r\in\{1,2,3\}.$ Accordingly to (1) the domain of the underlying trivariate subcopula is the set $\mbox{Dom}\,S^{(3)}=\{0,\theta_{1},1\}\times\{0,\theta_{2},1\}\times\{0,\theta_{3},1\}$ and:
$S^{(3)}(u,v,w)=\begin{cases}\,1&\mbox{ if }(u,v,w)=(1,1,1),\\ \,\theta_{1}&\mbox{ if }(u,v,w)=(\theta_{1},1,1),\\ \,\theta_{2}&\mbox{ if }(u,v,w)=(1,\theta_{2},1),\\ \,\theta_{3}&\mbox{ if }(u,v,w)=(1,1,\theta_{3}),\\ \,\theta_{12}&\mbox{ if }(u,v,w)=(\theta_{1},\theta_{2},1),\\ \,\theta_{13}&\mbox{ if }(u,v,w)=(\theta_{1},1,\theta_{3}),\\ \,\theta_{23}&\mbox{ if }(u,v,w)=(1,\theta_{2},\theta_{3}),\\ \,\theta_{123}&\mbox{ if }(u,v,w)=(\theta_{1},\theta_{2},\theta_{3}),\\ \,0&\mbox{ elsewhere.}\end{cases}$ (14)
@@
$\max\{\theta_{r}+\theta_{t}-1,0\}\,\leq\,\theta_{rt}\,\leq\,\min\{\theta_{r},\theta_{t}\}\,,\quad 1\leq r<t\leq 3\,.$ (15)
@@
but it will be discussed later that (16) is a necessary but could not be a sufficient condition for an admisible value for $\theta_{123},$ due to a compatibility problem, similarly as mentioned in the bivariate case. Since $\mbox{Ran}\,(X_{1},X_{2},X_{3})=\{(i,j,k):i,j,k\in\{0,1\}\}$ let:
$p_{ijk}:=\mathbb{P}(X_{1}=i,X_{2}=j,X_{3}=k)\,,\qquad i,j,k\in\{0,1\},$ (17)
@@
$\theta_{123}\,\leq\,\min\{\theta_{12},\theta_{13},\theta_{23}\}\,\leq\,\min\{\theta_{1},\theta_{2},\theta_{3}\}$ (27)
@@
Algorithm 1
  • Step 0:
    Choose $0\leq p_{000}\leq 1$ and set the trivariate dependence parameter $\theta_{123}=p_{000}\,.$
  • Step 1:
    Choose $0\leq p_{001}\leq 1,$ $0\leq p_{010}\leq 1,$ and $0\leq p_{100}\leq 1,$ such that $p_{000}+p_{001}+p_{010}+p_{100}\leq 1,$ and calculate the bivariate dependence parameters $\theta_{rt}$ accordingly to (29), (30), and (31).
  • Step 2:
    Choose $0\leq p_{011}\leq 1,$ $0\leq p_{101}\leq 1,$ and $0\leq p_{110}\leq 1,$ such that $p_{000}+p_{001}+p_{010}+p_{100}+p_{011}+p_{101}+p_{110}\leq 1,$ and calculate the univariate (marginal) parameters $\theta_{r}$ accordingly to (32), (33), and (34).
  • Step 3:
    Calculate $p_{111}=1-p_{000}-p_{001}-p_{010}-p_{011}-p_{100}-p_{101}-p_{110}\,.$
@@
where from (16) $-1\leq\mu_{123}\leq+1,$ and clearly $\mu_{123}=0$ if and only if the three random variables are independent.
@@
$\max\{0,2\theta^{2}-\theta\}\,\leq\,\theta_{123}\,\leq\,\min\{\theta^{2},1+\theta-2\theta^{2},1-3\theta+3\theta^{2}\}$ (36)
@@
$\theta_{123}\in\begin{cases}\,\left[0\,,\,\theta^{2}\right]&\text{ if }\theta\leq\frac{1}{2}\\ \,\left[2\theta^{2}-\theta\,,\,1-3\theta+3\theta^{2}\right]&\text{ if }\frac{1}{2}<\theta\leq\frac{4}{5}\\ \,\left[2\theta^{2}-\theta\,,\,1+\theta-2\theta^{2}\right]&\text{ if }\frac{4}{5}<\theta\leq\frac{1+\sqrt{5}}{4}\end{cases}$ (37)
@@
Figure 2: Compatible values $(\theta,\theta_{123})$ (red region)
in Example 1 for a Trivariate Bernoulli, with identical univariate marginals with parameter $1-\theta,$ pairwise independent, with trivariate dependence parameter $\theta_{123}.$
@@
$(0,0,0)$ $\theta_{123}$ 00 $1/8$ $1/4$
$(0,0,1)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 00
$(0,1,0)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 00
$(1,0,0)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 00
$(0,1,1)$ $\theta-2\theta^{2}+\theta_{123}$ 00 $1/8$ $1/4$
$(1,0,1)$ $\theta-2\theta^{2}+\theta_{123}$ 00 $1/8$ $1/4$
$(1,1,0)$ $\theta-2\theta^{2}+\theta_{123}$ 00 $1/8$ $1/4$
$(1,1,1)$ $1-3\theta+3\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 00
@@
  • Step 1)
    Choose $p_{000}\in[0,1]$ and set the trivariate dependence parameter $\tau=p_{000}.$
  • Step 2)
    Choose $p_{001},p_{010},p_{100}\in[0,1]$ such that:
    $p_{000}+p_{001}+p_{010}+p_{100}\leq 1$ (38)
@@
  • Step 3)
    Choose $p_{011},p_{101},p_{110}\in[0,1]$ such that:
@@
  • Step 4)
    Calculate $p_{111}=1-\tau-3(\beta-\tau)-3(\theta-2\beta+\tau)=1-\tau+3(\beta-\theta).$

### v2

@@ Abstract
By applying Sklar’s theorem to the Multivariate Bernoulli Distribution (MBD), this paper proposes a framework to decouple marginal distributions from the dependence structure, clarifying interactions among binary variables. Explicit formulas are derived under the MBD using subcopulas to introduce dependence measures for interactions of all orders, not just pairwise. A Bayesian inference approach is also applied to estimate the parameters of the MBD, offering practical tools for parameter estimation and dependence analysis in real-world applications. The results obtained contribute to the application of subcopulas of multivariate binary data, with real data examples.
keywords:
Sklar’s theorem, subcopula, dependence, Bernoulli, binary data
@@ 1 Introduction
A representation of the MBD that provides an alternative to the traditional log-linear model for binary variables was proposed by [1] using the concept of Kronecker product from matrix calculus, in terms of $2^{n}-1$ parameters, where $n$ is the dimension of a random vector with MBD, but where the dependency vector of parameters is, in fact, non standardized central moments. In [2] several flexible methods for simulating random binary sequences with fixed marginal distributions and specified degrees of association between the variables are discussed, but only pairwise dependencies are considered. An algorithm is derived in [3] for generating systems of correlated binary data, but allowing for the specification of just pairwise correlations within each system. A simple method to characterize multivariate Bernoulli variables with given means is investigated in [4] but only pairwise correlations are discussed.
Copulas, and more generally subcopulas, provide a straightforward approach to describing the dependence structure in multivariate settings by decoupling the marginal distributions from the joint distribution. Introduced by Sklar’s theorem [5], subcopulas extend copula theory to cases where the marginals are not necessarily continuous, making them particularly useful for multivariate discrete data.
In [6] it is discussed that many of the convenient properties of copulas do not carry over from the continuous to the discrete case, because the underlying copula is not unique, but according to [5] there is a unique subcopula, and that is the reason why this paper focuses on leveraging subcopulas to characterize the dependence within the MBD, offering a framework to better understand and quantify dependence of all orders (pairwise, three-wise, four-wise, etc.)
The present work introduces a formal subcopula-based characterization of the dependence structure within the MBD, and illustrates its usefulness to tackle the compatibility problem of building multivariate models in terms of lower dimension marginals.
@@ 2 Subcopulas and dependence
Sklar’s Theorem [5] proves that for any $n$-dimensional vector of random variables $(X_{1},\ldots,X_{n})$ there exists a functional link $S^{(n)}$ between its joint probability distribution $F_{X_{1},\ldots,X_{n}}(x_{1},\ldots,x_{n})=\mathbb{P}(X_{1}\leq x_{1},\ldots,X_{n}\leq x_{n})$ and its corresponding univariate marginal distributions $F_{X_{i}}(x_{i})=\mathbb{P}(X_{i}\leq x_{i}),$ $i\in\{1,\ldots,n\}:$
@@
for any $(x_{1},\ldots,x_{n})\in\overline{\mathbb{R}}^{n},$ and where the function $S^{(n)}:D_{1}\times\cdots\times D_{n}\rightarrow[0,1]$ is unique, with $D_{i}=\mbox{Ran}\,F_{X_{i}}=\{F_{X_{i}}(x):x\in\overline{\mathbb{R}}\}$ for $i\in\{1,\ldots,n\},$ which implies $\{0,1\}\subseteq D_{i}\subseteq[0,1].$ Such link $S^{(n)}$ is called a subcopula function, and in the particular case where every $D_{i}=[0,1]$ then it is called a copula function, which would be the case when all the random variables are continuous. If at least one of the random variables is not continuous then the domain of the underlying subcopula $S^{(n)}$ is a proper subset of the unit hypercube $0,1]^{n}.$ A formal definition for subcopula functions and their properties are discussed in detail in [8] and [9].
Recalling that the random variables $X_{1},\ldots,X_{n}$ are mutually independent if and only if their probability joint distribution is equal to the product of its univariate marginals, then as an immediate consequence of (1) the unique subcopula function under independence is:
@@
Moreover, as a consequence of the Fréchet-Hoeffding bounds for joint distributions (see [10] and [11]) and, applying (1), we may obtain bounds for any subcopula $S^{(n)}:$
@@
From any $n$-dimensional subcopula $S^{(n)}(u_{1},\ldots,u_{n})$ and $2\leq k<n$ it is possible to obtain $\binom{n}{k}$ $k$-dimensional marginal subcopulas $S^{(k)}(v_{1},\ldots,v_{k})$ defined in terms of $S^{(n)}(u_{1},\ldots,u_{n})$ where $k$ of its entries are equal to exactly one of each $\{v_{1},\ldots,v_{k}\}$ and the remaining nkn-k are equal to $1.$
Since the univariate marginal distributions $F_{X_{i}}(x_{i})=\mathbb{P}(X_{i}\leq x_{i})$ have no information about how each random variable $X_{i}$ interacts with others, then as a consequence of (1) all the information about the dependence among the random variables $X_{1},\ldots,X_{n}$ is contained in their unique underlying subcopula $S^{(n)},$ and therefore any attempt to measure degrees or intensity of dependence should extract information from $S^{(n)}.$
For the particular case of bivariate copulas (when both random variables are continuous) several copula-based measures have been studied, such as the concordance measures by Spearman [12] or Kendall [13], or dependence measures by Schweizer-Wolff [14] or Hoeffding [11].
But when at least one of the variables is discrete, as pointed out by [6], an identifiability issue arises, and therefore copula-based dependence or concordance measures cannot be adapted in this case since it is not possible to report a unique value. This motivated the work by [15] where for the case of proper subcopulas (that are not copulas) a subcopula dependence measure has been proposed. For the case of a vector $(X,Y)$ of arbitrary type random variables with underlying and unique subcopula $S$ it is defined:
$d(S):=\sup_{\text{Dom}\,S}\{S-\Pi_{S}\}-\sup_{\text{Dom}\,S}\{\Pi_{S}-S\}\,.$ (4)
where for any $(u,v)\in\text{Dom}\,S=\text{Ran}\,F_{X}\times\text{Ran}\,F_{Y}$ the following subcopulas are defined as $W_{S}(u,v):=\max\{u+v-1,0\},$ $M_{S}(u,v):=\min\{u,v\},$ and $\Pi_{S}(u,v)=uv,$ and then:
[…]
where $d(M_{S})$ and $d(W_{S})$ is formula (4) applied to $M_{S}$ and $W_{S},$ respectively. In (4) it is calculated the largest difference between the unique underlying subcopula $S$ and the subcopula $\Pi_{S}$ that represents independence, and this value is used in (5) to calculate a proportion with respect to the largest possible differences, according to the Frécht-Hoeffding bounds for subcopulas. In fact, (4) and (5) may be similarly applied to any subcopula in higher dimensions, using (2) and (3).
@@ 3 Bivariate Bernoulli
Let $(X_{1},X_{2})$ be a vector of Bernoulli random variables with (univariate marginal) parameters $1-\theta_{r}$ where $0<\theta_{r}<1,$ that is $\mathbb{P}(X_{r}=0)=\theta_{r}$ and $\mathbb{P}(X_{r}=1)=1-\theta_{r}$ for $r\in\{1,2\}.$ Then the univariate marginal distribution functions are given by $F_{r}(x)=\mathbb{P}(X_{r}\leq x)=\theta_{r}\mathbb{I}_{\{0\,\leq\,x\,<\,1\}}+\mathbb{I}_{\{x\,\geq\,1\}},$ where $\mathbb{I}_{A}$ stands for the indicator (or characteristic) function of the set $A,$ and therefore $\mbox{Ran}\,F_{r}=\{0,\theta_{r},1\}.$
According to (1) the domain of the underlying subcopula in this case is the set $\mbox{Dom}\,S^{(2)}=\{0,\theta_{1},1\}\times\{0,\theta_{2},1\}.$ From the formal definition of bivariate subcopula (see [9], for example) it always satisfies $S^{(2)}(u,0)=0=S^{(2)}(0,v),$ $S^{(2)}(u,1)=u,$ and $S^{(2)}(1,v)=v,$ and therefore:
@@
Formulas (8) allow us to fully characterize the family of all bivariate Bernoulli distributions in terms of the marginal parameters $0<\theta_{1}<1$ and $0<\theta_{2}<1,$ and a bivariate dependence parameter $\theta_{12}$ that must satisfy $\max\{\theta_{1}+\theta_{2}-1,0\}\leq\theta_{12}\leq\min\{\theta_{1},\theta_{2}\},$ that we may denote as $\mathcal{B}_{2}(\theta_{1},\theta_{2},\theta_{12}).$ The subcopula based dependence measure proposed by [15] for $\mathcal{B}_{2}(\theta_{1},\theta_{2},\theta_{12})$ becomes:
@@
$(p_{00},p_{01},p_{10},p_{11})=(\theta_{1}\theta_{2},\theta_{1}(1-\theta_{2}),(1-\theta_{1})\theta_{2},(1-\theta_{1})(1-\theta_{2}))\,$,
@@
$(p_{00},p_{01},p_{10},p_{11})=(\theta_{1},0,\theta_{2}-\theta_{1},1-\theta_{2})\,.$
@@
[…]
@@
In summary, formulas (8) allow to define the family $\mathcal{B}_{2}(\theta_{1},\theta_{2},\theta_{12})$ of all bivariate distributions with Bernoulli univariate marginals, which is in fact also known as a Fréchet class of the type $\mathcal{F}(F_{1},F_{2}),$ see for example [16]. A Fréchet class is a set of all multivariate distribution functions that share the same given marginal distributions. Formally, let $F_{1},F_{2},\ldots,F_{d}$ be univariate cummulative distribution functions. The Fréchet class associated with $(F_{1},\ldots,F_{d})$ is defined as:
$\mathcal{F}(F_{1},\ldots,F_{d}):=\{H:H\text{ is a $d$-dimensional distribution with marginals }F_{1},\ldots,F_{d}\}$
The bivariate joint probability mass function $p_{ij}=\mathbb{P}(X_{1}=i,X_{2}=j)$ for $i$ and $j$ in $\{0,1\}$ can be specified in two ways:
@@
    choosing freely the two marginal parameters $0<\theta_{1}<1$ and $0<\theta_{2}<1$ and then choosing $\theta_{12}$ such that (7) holds, or
@@
    choosing freely the dependence parameter $0<\theta_{12}<1$ and then choosing the marginal parameters $0<\theta_{1}<1$ and $0<\theta_{2}<1$ such that (7) holds. Equivalently, instead of choosing the dependence parameter value, we may choose a given value of the dependence measure defined by [15] and then obtain the value for $\theta_{12}$ from (9).
Behind the above cases a) and b) is what is known as a compatibility problem, in this particular case for the Fréchet class $\mathcal{F}(F_{1},F_{2}):$ even though the three parameters are the probabilities $\theta_{1}=\mathbb{P}(X_{1}=0),$ $\theta_{2}=\mathbb{P}(X_{2}=0),$ and $\theta_{12}=\mathbb{P}(X_{1}=0,X_{2}=0),$ not all values in the $[0,1]$ interval are simultaneously admissible (or compatible) for such parameters, since (7) must hold, see Figure 1:
  • a)
    Figure 1 Left: Specific values are chosen for $\theta_{1}=\theta_{1}^{*}$ and $\theta_{2}=\theta_{2}^{*}$ and from (7) all the admissible values for $\theta_{12}\in\big{[}\max\{\theta_{1}^{*}+\theta_{2}^{*}-1,0\}\,,\,\min\{\theta_{1}^{*},\theta_{2}^{*}\}\big{]}.$ If the chosen point $(\theta_{1}^{*},\theta_{2}^{*})$ is below the solid gray identity line $\theta_{2}=\theta_{1}$ then $\min\{\theta_{1}^{*},\theta_{2}^{*}\}=\theta_{1}^{*},$ which is projected to the horizontal axis through such identity line (horizontally through black solid line, then vertically through a violet solid line to point $(0,\theta_{1}^{*})$), to get the upper bound for $\theta_{12}.$ For the lower bound, since $\theta_{1}+\theta_{2}-1\geq 0$ if and only if $\theta_{2}\geq 1-\theta_{1},$ then if the chosen point $(\theta_{1}^{*},\theta_{2}^{*})$ is above the solid gray line $\theta_{2}=1-\theta_{1},$ we may build a solid black line $\theta_{2}=-\theta_{1}+(\theta_{1}^{*}+\theta_{2}^{*})$ that includes the point $(\theta_{1}^{*},\theta_{2}^{*})$ and that is parallel to the solid gray line $\theta_{2}=1-\theta_{1},$ so that $\theta_{2}=1$ if and only if $\theta_{1}=\theta_{1}^{*}+\theta_{2}^{*}-1,$ which is the corresponding lower bound for $\theta_{12},$ and so the point $(\theta_{1}^{*}+\theta_{2}^{*}-1,1)$ is projected vertically down to $(0,\theta_{1}^{*}+\theta_{2}^{*}-1).$ So the red thick line represents the interval $\theta_{1}^{*}+\theta_{2}^{*}-1\,,\,\theta_{1}^{*}\,]$ of admissible values for $\theta_{12}$ in this graphical example, given chosen values $\theta_{1}=\theta_{1}^{*}$ and $\theta_{2}=\theta_{2}^{*}.$
  • b)
    Figure 1 Right: In this case what is chosen is a specific value $\theta_{12}=\theta_{12}^{*}$ and now will characterize the set of all points $(\theta_{1},\theta_{2})$ that are compatible with such fixed value $\theta_{12}^{*}$ according to (7), that is to solve for $\theta_{1}$ and $\theta_{2}$ these two inequalities:
    $\max\{\theta_{1}+\theta_{2}-1,0\}$ $\leq$ $\theta_{12}^{*}\,$,
    $\min\{\theta_{1},\theta_{2}\}$ $\geq$ $\theta_{12}^{*}\,$,
    which in turn are equivalent to the following inequalities:
    $\theta_{2}$ $\leq$ $\min\{1,-\theta_{1}+\theta_{12}^{*}+1\}\,$,
[…]
    where these last two inequalities represent the red triangle region of admissible values for $(\theta_{1},\theta_{2})$ given a fixed value $\theta_{12}=\theta_{12}^{*}.$
@@ 4 Trivariate Bernoulli
Let $(X_{1},X_{2},X_{3})$ be a vector of Bernoulli random variables with (univariate marginal) parameters $1-\theta_{r}$ where $0<\theta_{r}<1$ as explained in the previous section, but now $r\in\{1,2,3\}.$ According to (1) the domain of the underlying trivariate subcopula is the set $\mbox{Dom}\,S^{(3)}=\{0,\theta_{1},1\}\times\{0,\theta_{2},1\}\times\{0,\theta_{3},1\}$ and:
[…]
@@
$\max\{\theta_{r}+\theta_{t}-1,0\}\,\leq\,\theta_{rt}\,\leq\,\min\{\theta_{r},\theta_{t}\}\,,\quad 1\leq r<t\leq 3\,.$
@@
but it will be discussed later that (11) is a necessary but could not be a sufficient condition for an admissible value for $\theta_{123},$ due to a compatibility problem, similarly as mentioned in the bivariate case. Since $\mbox{Ran}\,(X_{1},X_{2},X_{3})=\{(i,j,k):i,j,k\in\{0,1\}\}$ let:
$p_{ijk}:=\mathbb{P}(X_{1}=i,X_{2}=j,X_{3}=k)\,,\qquad i,j,k\in\{0,1\},$
@@
$\theta_{123}\,\leq\,\min\{\theta_{12},\theta_{13},\theta_{23}\}\,\leq\,\min\{\theta_{1},\theta_{2},\theta_{3}\}$
@@
Parameter Selection Procedure 1
  • Step 1: Choose $0\leq p_{000}\leq 1$ and set the trivariate dependence parameter $\theta_{123}=p_{000}\,.$
  • Step 2: Choose $0\leq p_{001}\leq 1,$ $0\leq p_{010}\leq 1,$ and $0\leq p_{100}\leq 1,$ such that $p_{000}+p_{001}+p_{010}+p_{100}\leq 1,$ and calculate the bivariate dependence parameters $\theta_{rt}$ according to (21), (22), and (23).
  • Step 3: Choose $0\leq p_{011}\leq 1,$ $0\leq p_{101}\leq 1,$ and $0\leq p_{110}\leq 1,$ such that $p_{000}+p_{001}+p_{010}+p_{100}+p_{011}+p_{101}+p_{110}\leq 1,$ and calculate the univariate (marginal) parameters $\theta_{r}$ according to (24), (25), and (26).
  • Step 4: Calculate $p_{111}=1-p_{000}-p_{001}-p_{010}-p_{011}-p_{100}-p_{101}-p_{110}\,.$
@@
where from (11) $1\leq\mu_{123}\leq+1,$ and clearly $\mu_{123}=0$ if and only if the three random variables are independent.
@@
$\max\{0,2\theta^{2}-\theta\}\,\leq\,\theta_{123}\,\leq\,\min\{\theta^{2},1+\theta-2\theta^{2},1-3\theta+3\theta^{2}\}$
@@
[…]
@@
Figure 2: Compatible values $(\theta,\theta_{123})$ (red region) in Example 1 for a Trivariate Bernoulli, with identical univariate marginals with parameter $1-\theta,$ pairwise independent, with trivariate dependence parameter $\theta_{123}.$
@@
$(0,0,0)$ $\theta_{123}$ 0 $1/8$ $1/4$
$(0,0,1)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 0
$(0,1,0)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 0
$(1,0,0)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 0
$(0,1,1)$ $\theta-2\theta^{2}+\theta_{123}$ 0 $1/8$ $1/4$
$(1,0,1)$ $\theta-2\theta^{2}+\theta_{123}$ 0 $1/8$ $1/4$
$(1,1,0)$ $\theta-2\theta^{2}+\theta_{123}$ 0 $1/8$ $1/4$
$(1,1,1)$ $1-3\theta+3\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 0
@@
  • Step 1: Choose $p_{000}\in[0,1]$ and set the trivariate dependence parameter $\tau=p_{000}.$
  • Step 2: Choose $p_{001},p_{010},p_{100}\in[0,1]$ such that:
    $p_{000}+p_{001}+p_{010}+p_{100}\leq 1$
@@
  • Step 3: Choose $p_{011},p_{101},p_{110}\in[0,1]$ such that:
@@
  • Step 4: Calculate $p_{111}=1-\tau-3(\beta-\tau)-3(\theta-2\beta+\tau)=1-\tau+3(\beta-\theta).$